# TalkPlayData 2: An Agentic Synthetic Data Pipeline for Multimodal Conversational Music Recommendation

### v1

@@ 5 Conclusion
In this paper, we introduced TalkPlayData 2, a new multimodal dataset for conversational recommendation systems. In the data generation pipeline, separate LLM calls are first made to create a listener profile and a conversation goal for each conversation. Using them as a condition, two separate LLMs talk to each other under the role of a music listener and a music recommendation system. The conversation is conducted for 8 turns, and the data is collected as a conversation. Notably, all the LLMs are multimodal, enabling to generate conversation with multimodal aspects of music being considered. The LLMs have access to different subsets of the information, a design choice that is highly similar to the real-world conversational recommendation systems. In the evaluation, we conducted a LLM-as-a-judge evaluation as well as a human evaluation, which shows that TalkPlayData 2 is a promising dataset for training and evaluating conversational music recommendation systems.
@@ B.2 Conversation Goal LLM
Output (YAML block of Conversation Goal, some are omitted for brevity) category_code: [alphabetical topic code among A-K] specificity_code: [one of $L$, HL, LH, $H$] target_turn_count: [1-8 based on specificity code] listener_goal: [customized goal description for the tracks] listener_expertise: [description of the listener expertise] initial_query_example_1: [plausible initial query example 1]

### v4

@@ 6 Conclusion
In this paper, we introduced TalkPlayData 2, a new multimodal dataset for conversational recommendation systems. In the data generation pipeline, separate LLM calls are first made to create a listener profile and a conversation goal for each conversation. Using them as a condition, two separate LLMs talk to each other under the role of a music listener and a music recommendation system. The conversation is conducted for 8 turns, and the data is collected as a conversation. Notably, all the LLMs are multimodal, enabling to generate conversation with multimodal aspects of music being considered. The LLMs have access to different subsets of the information, a design choice that is highly similar to the real-world conversational recommendation systems. In the evaluation, we conducted an LLM-as-a-judge evaluation as well as a human evaluation, which shows that TalkPlayData 2 is a promising dataset for training and evaluating conversational music recommendation systems.
@@ B.2 Conversation Goal LLM
Output (YAML block of Conversation Goal, some are omitted for brevity) category_code: [alphabetical topic code among A-K] specificity_code: [one of LL, HL, LH, $H$] target_turn_count: [1-8 based on specificity code] listener_goal: [customized goal description for the tracks] listener_expertise: [description of the listener expertise] initial_query_example_1: [plausible initial query example 1]